OpenStudy (anonymous):

hi..i'm trying to help my daughter find the least common denominator I understand that part; she is finding the multiples it says to write an equivalent fraction for each fraction how do I do that could you please help me with this one? 1/8 and 1/4

OpenStudy (anonymous):

1/2 not sure

OpenStudy (radar):

To write an equivalent fraction you may multiply or divide both the numerator and denominator by the same number. For example 1/8 multiply top and bottom by 2 getting 2/16 that being equivalent to 1/8 1/4 multiply top and bottom by 3 getting 3/12 that being equivalent to 1/4 6/12 divide top and bottom by 6 getting 1/2 that being equivalent to 6/12 do you get the drift, the key is to do the same to both top and bottom so you dont change the value of the fraction or making in unequivalent.

OpenStudy (anonymous):

hi radar...okay..so we need to find an equivalent fraction for each fraction, correct? then why would we need to find the least common multiple as it wants us to do? doesn't it want us to make the two fractions equivalent?

OpenStudy (anonymous):

the directions read : Find the LCD of the pair of fractions. then use the LCD to write an equivalent fraction for each fraction

OpenStudy (radar):

To add or subtract fractions the denominator of the fractions to be added (or subtracted) must be the same or common. In the example to add your 1/8 to the 1/4 there has to be a conversion so that both the denominator are the same. this is done by changeing the fraction to an equivalennt fraction with the same denominators. The LCD for 1/4 and 1/8 is 8 1/8 you leave alone, but the 1/4 has to become an equivalent fraction with a denominator 8. You do this by multiplying the 4 (denominator) by 2. BUT, but, but you must multiply the numerator by 2 also. Now the fraction appears as 2/8 (but it has not changed in value just appearance 2/8 is equivalent to 1/4 Now we can add. 2/8 + 1/8 = 3/8

OpenStudy (anonymous):

okay..that's what I thought; it doesn't want them to add the fractions together just find an equivalent

OpenStudy (radar):

Oh OK, but when you do add the denominators should be the same then the sum is simply the sum of the numerators over the common denominator.

OpenStudy (radar):

1/8, 2/16, 3/24, 5/40 are all equivalent.

OpenStudy (anonymous):

okay..so with 2/5 the equivalent fraction would be 4/10, right? 7/15 would 14/30?

OpenStudy (anonymous):

35/100 be equivalent to 70/200 9/20 =18/40 correct?

OpenStudy (radar):

2/5 2 is the numerator, 5 is the denominator multiply the numerator and the denominator by the same and you will come up with an equivalent fraction. your 2/5 could ber4/10 and hyour 7/15 could be 14/30 yes

OpenStudy (radar):

Yes you are on to it.

OpenStudy (anonymous):

thank you so much..i appreciate your help I will definitely be back to this site for more help!!!

OpenStudy (radar):

O.K., you have been giving examples where you were using 2 as a multiplier but 3 or 4 or 5 could also been used.

OpenStudy (radar):

Just be sure to multiply both the top and bottom whatever number you should choose.

OpenStudy (radar):

Good luck, and bless you for helping your daughter.